### v1

# Leibniz’s contested infinitesimals: Further depictions

Abstract.
We contribute to the lively debate in current scholarship on the Leibnizian calculus. In a recent text, Arthur and Rabouin argue that non-Archimedean continua are incompatible with Leibniz’s concepts of number, quantity and magnitude.
They allege that Leibniz viewed infinitesimals as contradictory, and claim to deduce such a conclusion from an analysis of the Leibnizian definition of quantity. However, their argument is marred by numerous errors, deliberate omissions, and misrepresentations, stemming in a number of cases from flawed analyses in their earlier publications.
We defend the thesis, traceable to the classic study by Henk Bos, that Leibniz used genuine infinitesimals, which he viewed as fictional mathematical entities (and not merely shorthand for talk about more ordinary quantities) on par with negatives and imaginaries.
Key words and phrases:
Infinitesimals; Archimedean property; accidental impossibility; absolute impossibility; infinita terminata; Leibniz
2020 Mathematics Subject Classification:
Primary 01A45, 01A61 Secondary 01A85, 01A90, 26E35
It is true that, in my view, infinites are not wholes, and the infinitely small are not magnitudes. My metaphysics banishes them from its territories, and I don’t accord them residence except in the imaginary spaces of the geometric calculus, where such notions are only as proper as the roots that one calls imaginary. The role I played in making this calculus of infinitesimals appreciated does not make me sufficiently enamored to push them beyond common sense.
–Leibniz to Fontenelle, 9 september 1705 (translation ours).
1. Introduction
In 2022, Arthur, Rabouin, and others published a brief two-page broadside [2], essentially free of serious content, against the work of scholars who challenge the currently received dogma on the Leibnizian calculus specifically and other historical infinitesimalists more generally. A detailed rebuttal appeared in Bair et al. [9] (a brief response appeared in [10]). Apparently feeling that the matter requires a more thorough treatment, Arthur and Rabouin recently published a 17-page follow-up [6], containing a critique of what they describe as “the vast program launched in the last ten years or so … to reinterpret the history of infinitesimal techniques, …” [^1] Arthur and Rabouin [6, note 2]. We are grateful to Arthur and Rabouin for giving us an opportunity to develop a detailed assessment of their position. As we analyze below, their new article contains significant errors and misinterpretations of Leibniz’s texts.
Richard Arthur and David Rabouin (henceforth AR) announce that they are pursuing a new strategy so as to tackle the question of the existence of infinitely small quantities in the Leibnizian calculus. Such a strategy involves “the study of his conception of quantity, number or magnitude.” [^2] Op. cit., p. 27. They claim to have shown that Leibnizian views on what a quantity, a magnitude or a number have to be, “are not compatible with the existence of non-Archimedean instances thereof.” [^3] Op. cit., p. 28. They claim further that the ongoing debates about whether nonstandard analysis (NSA) is an appropriate framework for reinterpreting the Leibnizian calculus rest on a series of misunderstandings, primarily on a “systematic conflation of the question of use with that of existence of infinitely small quantities.” [^4] Op. cit., p. 27. (on use and existence see Section 8).
In reality, AR themselves succumb to some significant conflations. We point out three major errors in their argumentation:
  1. (1)
    the conflation of existence in nature and existence in mathematics,
  2. (2)
    the conflation of infinite wholes and bounded infinities (infinita terminata),
  3. (3)
    the conflation of homogeneity and comparability.
In Sections 2 through 4, we address these errors in more detail. Additional misrepresentations and misconceptions in AR’s text are analyzed in Sections 5 through 8. Section 9 contrasts the position of AR and that of Henk Bos and other scholars. Section 10 chronicles Rabouin’s about-face concerning the presence of “limit-processes” in the Leibnizian calculus.
2. Existence in nature and existence in mathematics
AR’s main argument for the incompatibility of Leibniz’s views with non-Archimedean continua is that Leibniz described his inassignables as impossible entities, as contradictory and as absurd. According to AR, Leibniz proved the absurdity of infinitely small lines. They claim further: “we can exhibit a text in which Leibniz finally elaborated this proof in full detail: the notes he took on Froidmont.” [^5] Op. cit., p. 38. AR go on to present the following Leibnizian passage:
“Therefore at some time the ratio of what is traversed to the unassignable whole must necessarily begin to be assignable, but such a point or instant in which that transition comes about is impossible. Therefore such a motion, and so such a line, must also be held to be impossible, and there is no infinite yet bounded straight line. Whence it also follows that infinitely small straight lines are chimeras, although nevertheless useful, like imaginary roots …” [^6] Leibniz as translated in op. cit., p. 39; emphasis added.
Note that Leibniz speaks of motion (a physical notion). In fact, a few lines earlier AR acknowledge that
The general context is a part of the notes in which Leibniz goes back to the possibility of interpreting physical conatus as infinitely small lines traversed in a moment. [^7] Op. cit., pp. 38–39.
Evidently, the issue is the question of the existence of infinitesimals in the physical realm. Failing to distinguish the latter sufficiently from the mathematical realm, [^8] Arthur has a history of conflating these two Leibnizian realms by applying actuality to the mathematical realm, as we discuss in Section 4.3. AR argue as follows:
There is no ambiguity about the conclusion here: Leibniz considered such a reasoning as establishing the impossibility of infinite bounded lines, so that they should be ranked with infinitely small lines in the store of chimerical mathematical entities. But the most important fact is to note that from 1676 to the reflections on Froidmont, we see Leibniz considering infinitesimal lines as absurd, even when not taken as genuine magnitudes. [^9] Arthur and Rabouin [6, p. 39]; emphasis added.
So far, infinitesimals are impossible, chimerical, and absurd. But then AR jump to the following conclusion:
As we have established, he attempted regularly to prove that they involve a contradiction. [^10] Ibid.; emphasis added.
Did Leibniz attempt to prove that infinitesimals involve a contradiction, as AR claim? As AR correctly point out some pages earlier,
According to Leibniz, mathematical existence can be equated with non-contradiction. [^11] Op. cit., p. 27, note 6.
If so, how could Leibniz have used infinitesimals if they are contradictory according to AR? We will examine the matter in Section 2.1.
2.1. Accidental and absolute impossibility
What AR appear to ignore is the fact that Leibniz distinguished between existence in the phenomenal realm (i.e., physical realm) and mathematical existence (which also includes fictional entities).
For Leibniz, fictional did not necessarily mean contradictory. Crucial here is Leibniz’s distinction between absolute impossibility and accidental impossibility. Already in the Confessio Philosophi, [^12] Leibniz [40], 1672/3, p. 128. Leibniz speaks of the concept of “impossible by accident” and contrasts it with “absolute impossibility” i.e., contradiction. He gives the examples of a species with an odd number of feet, and an immortal mindless creature, which are, according to him, harmoniae rerum adversa i.e., “contrary to the harmony of things,” [^13] Leibniz as translated by Sleigh in [48, p. 57]. but not necessarily contradictory.
In his 1683 text Elementa nova matheseos universalis, Leibniz explains that some mathematical operations cannot be performed in actuality, but nonetheless one can exhibit
“a construction in our characters” (in nostris characteribus), [^14] Leibniz [41, p. 520].
meaning that one can carry out a formal calculation, such as those with imaginary roots, regardless of whether the mathematical notions involved idealize anything in nature. Leibniz goes on to discuss in detail the cases of imaginary roots and infinitesimals, which to him constitute examples of accidental impossibility. [^15] For a detailed textual analysis, see [33, Section 5.3]. The distinction was also analyzed by Eklund [18].
Something contradictory is absolutely impossible and can neither occur in nature nor be allowed in mathematics. It may be absurd to assume the existence of negatives, imaginaries, or infinitesimal quantities in nature. [^16] AR’s article contains several relevant passages from Leibniz: “whether the nature of things tolerates quantities of this kind is for the metaphysician to discuss; it is enough for the geometrician to demonstrate what follows from supposing them” [6, p. 35]; “I would indeed admit these infinitely small spaces and times in geometry for the sake of discovery, even if they are imaginary. But I doubt whether they can be admitted in nature” (op. cit., p. 37). However, such entities are not contradictory, but useful for mathematics. Leibniz therefore described them as well-founded fictions. Such a viewpoint involves no conflation of use and existence of infinitesimals; see further in Section 8. AR’s conclusion that
from 1676 to the reflections on Froidmont, we see Leibniz considering infinitesimal lines as absurd, …  he attempted regularly to prove that they involve a contradiction [^17] Arthur and Rabouin [6, p. 39]; emphasis added.
is a non-sequitur, because infinitesimals can be “absurd” without being contradictory. AR claim further that
Leibniz’s metaphysics claims that only simple substances should be called ‘real’ stricto sensu and it is not easy to understand in this framework the distinction between fictitious and real entities in mathematics. [^18] Op. cit., note 46.
In fact, AR exaggerate the difficulty involved. Certainly, Leibniz considered only entities in a metaphysical realm as truly real. But as far as mathematics is concerned, ordinary mathematical entities are those that were familiar to 17th century scholars from their study of the Greeks. Meanwhile, entities of a new type: imaginaries, negatives, and infinitesimals [^19] Leibniz specifically compares the status of negatives and infinitesimals: “[J]ust as I deny the reality of a ratio one of whose terms is a quantity less than nothing, I also deny that properly speaking there exists an infinite or infinitely small number, or an infinite or infinitely small line, …” This passage is translated in [6, p. 30] by AR, who fail to examine its implications. would be fictions; see [62] for a fuller discussion.
2.2. An epigraph and an artful ellipsis
AR’s epigraph furnishes an interesting case study. They quote Leibniz as follows:
“Il est vrai que chez moi les infinis ne sont pas des touts et les infiniment petits ne sont pas des grandeurs. (…) [^20] The ellipsis is in the original. La part que j’ai eue à faire valoir ce calcul des infinitésimales ne me rend pas assez amoureux pour les pousser au-delà du bon sens.” (Leibniz à Fontenelle, 9 septembre 1705) [^21] Leibniz as quoted by Arthur and Rabouin [6, Epigraph]. A translation of the full passage appears in note 22.
We note the ellipsis in the middle of the quoted passage. The passage as presented makes it appear as though Leibniz is telling Fontenelle that genuine infinitesimals would go beyond common sense (“bon sens”). Was that truly Leibniz’s intention? A fuller quotation is readily found in an article by Bella. Here is the full passage, as quoted by Bella:
“Il est vrai que chez moi les infinis ne sont pas des touts et les infiniment petits ne sont pas des grandeurs. Ma métaphysique les bannit de ses terres, et je ne leur donne retrait que dans les espaces imaginaires du calcul géométrique, ou ces notions ne sont de mise que comme les racines qu’on appelle imaginaires. La part que j’ai eue à faire valoir ce calcul des infinitésimales ne me rend pas assez amoureux pour les pousser au-delà du bon sens.” [^22] Leibniz as cited in Bella [13, notes 77, 86]; emphasis added. Translation: “It is true that, in my view, infinites are not wholes, and the infinitely small are not magnitudes. My metaphysics banishes them from its territories, and I don’t accord them residence except in the imaginary spaces of the geometric calculus, where such notions are only as proper as the roots that one calls imaginary. The role I played in making this calculus of infinitesimals appreciated does not make me sufficiently enamored to push them beyond common sense.”
It emerges that, contrary to the picture AR seek to paint, the meaning of the Leibnizian passage is as follows:
  1. (1)
    it is Leibniz’s metaphysics that expels infinitesimals from “ses terres”;
  2. (2)
    Leibniz nonetheless leaves room for infinitesimals in the imaginary spaces of geometric calculation;
  3. (3)
    he compares infinitesimals to imaginary roots.
Accordingly, genuine infinitesimals are fine in mathematics just as imaginary roots are, but insisting on their existence beyond the realm of mathematics would push them beyond common sense, according to Leibniz. By suppressing the crucial middle sentence, AR changed the apparent meaning of the Leibnizian passage. The letter to Fontenelle proves precisely the opposite of what AR make it out to say by means of a skillful use of an ellipsis. [^23] Earlier skillful uses of the truncation technique to score an apparently syncategorematic point in the writings of Ferraro [22, p. 29], Goldenbaum [24, p. 76], Ishiguro [28], and Tho [61, p. 71] were analyzed in Bascelli et al. [11].
3. Infinite wholes and bounded infinities
Most commentators agree that Leibniz rejected infinite wholes as contradictory. From this, AR seek to deduce that he therefore also rejected the notion of a bounded infinite line (infinitum terminatum), [^24] The notion of infinitum terminatum is analyzed in Section 7. because the latter would consist of an infinite number of finite segments and such a “number” would have to be an infinite whole.
There is indeed a subtle difference between an infinite multitude of things (e.g., segments) as a whole, and a bounded but infinitely extended line consisting of an infinite number of segments. It is plain that infinite multitude was not a number for Leibniz, but the infinite extension of these segments was a tool useful in geometry. Leibniz distinguished between bounded infinity and infinite wholes, and indeed contrasted them. [^25] Thus, referring to infinita terminata as simply ‘infinita’, Leibniz writes: “quantitas interminata differt ab infinita” [51] A.VII. 6. 549. As noted by Knobloch, the distinction between infinita terminata and infinita interminata was elaborated in Proposition 11 of the treatise De Quadratura Arithmetica (DQA):
[Leibniz] distinguished between two infinites, the bounded infinite straight line, the recta infinita terminata, and the unbounded infinite straight line, the recta infinita interminata. He investigated this distinction in several studies from the year 1676. Only the first kind of straight lines can be used in mathematics, as he underlined in his proof of theorem 11 [i.e., Propositio XI]. [^26] Knobloch [36, p. 97].
Leibniz mentioned the distinction in a 29 july 1698 letter to Bernoulli. Here Leibniz analyzes a geometric problem involving unbounded infinite areas and states an apparent paradox: “Therefore the two infinite spaces are equal, and the part is equal to the whole, which is impossible.”272727“Ergo haec duo spatia infinita sibi sunt aequalia, pars toti, quod est absurdum” (Leibniz [45, p. 523]). To resolve the paradox (i.e., the clash with the part-whole principle), Leibniz exploits the notion of bounded infinity. Far from creating paradox, bounded infinities serve to resolve one. [^28] For a detailed textual analysis, see [35, Section 1]. Further details on infinita terminata appear in Section 7.
3.1. Whose wholes?
In their Conclusion, AR claim to “have shown that a quantity or magnitude is always defined as a ‘whole’ endowed with a certain number of (homogeneous) parts” [^29] Arthur and Rabouin [6, p. 41]. and go on to allege that since Leibniz rejected infinite wholes, he would necessarily also reject infinite quantity. How did they show this exactly? Their section 3 entitled “Number and quantity according to Leibniz” quotes Leibniz as follows:
“Quantity is that which belongs to a thing insofar as it has all its parts, or, on account of which it is said to be equal to, greater than, or less than, another thing (with each homogeneous with the other), or can be compared with it.” [^30] Leibniz as translated in Arthur and Rabouin [6, p. 29].
Note that ‘wholes’ are not mentioned in Leibniz’s passage. Immediately following this quote, AR proceed to paraphrase Leibniz:
According to this definition, a quantity is a whole constituted of parts and only things to which the part-whole axiom can be applied can count as quantities stricto sensu. [^31] Ibid.; emphasis on ‘whole’ added.
Note that the term whole only appears in AR’s paraphrase. Leibniz does not use the term in his definition of quantity. A similar paraphrase occurs when AR present a Leibnizian definition of magnitude:
Leibniz always defines continuous magnitudes, exactly like quantities, in terms of whole and parts: “Magnitude is that which is designated by a number of congruent parts”. If infinite wholes are contradictory, infinite magnitudes are contradictory notions from their very definition, and so are infinitely small parts of any finite magnitude. [^32] Arthur and Rabouin [6, p. 30]; emphasis added.
In this passage as well, the term whole is mentioned not by Leibniz but by AR, whose “contradictory” conclusion therefore depends on their paraphrase.
It emerges that AR have shown only that their paraphrases of the Leibnizian definitions of quantity and magnitude may lead to a rejection of infinite instances thereof, rather than the definitions as they appear in Leibniz.
3.2. Do infinite numbers require infinite sets?
AR claim that
Leibniz’s views about quantities, numbers and magnitude are not consistent with the NSA interpretation. This inconsistency takes the simplest form: the very notion of “quantity” in Leibniz, be they numbers or magnitudes, is not compatible with the fact of being “non-Archimedean”. [^33] Op. cit., p. 28; emphasis added.
Such an alleged ‘incompatibility’ is due to AR’s belief that an infinite quantity would necessarily be tied up with infinite sets as per AR (see Section 3.1), and therefore infinite quantities must be as contradictory as infinite sets in Leibniz’s world. AR detail their objections as follows:
There are several versions of nonstandard analyses [sic] depending on the kind of logical tools which are employed in order to introduce non-Archimedean quantities in mathematical analysis. Most of them, however, suppose strong set theoretic commitments like the existence of ultrapowers or conservative extensions of ZFC (as in various versions of Internal Set Theory). These commitments are not compatible with Leibniz’s conviction that any infinite totality is a contradictory notion … [^34] Op. cit., note 5; emphasis added.
Here AR speak of ‘strong set theoretic commitments’, ‘existence of ultrapowers’, and ‘ZFC’. The passage makes it appear as if infinite numbers in modern mathematics necessarily require ‘infinite totalities’ as well as the axiom of choice (part of ZFC), and are thus inapplicable to interpreting Leibniz, who rejected infinite totalities. But AR’s modern set-theoric excursion misses the point, as we explain in Section 3.3.
3.3. Conservativity results by Friedman and others
In the late 1960s, Harvey Friedman proved the conservativity of not only PA over PA, but the more powerful theory, namely PA plus the induction scheme referred to as ${}^{*}\Pi_{\infty}\mathsf{-Induction}$. Friedman’s work was mentioned by Kreisel. [^35] See Kreisel [38, p. 94]. See also Feferman [21, p. 967] (second theorem in Section 8.7) and Kossak [37]. Friedman’s result was followed up by Keisler et al. [26]. Thus, Robinson’s question concerning axiomatisations was answered by Friedman and led to a mature body of work. [^36] See Enayat [19] for further details.
Here PA denotes Peano Arithmetic, an axiomatic system that does not prove the existence of an infinite set. The extension PA incorporates the standardness predicate (as well as other axioms elaborated there) and therefore postulates the existence of nonstandard integers. By conservativity, PA does not prove the existence of an infinite set, either. Therefore the existence of an infinite – or more precisely, unlimited – integer does not require the existence of infinite sets.
To emphasize the point, one could refer to infinite integers in nonstandard models of arithmetic as ringinals, because they are elements of a (semi)ring, to contrast them with infinite numbers of Cantorian set theory such as ordinals or cardinals (which do not satisfy the properties of a ring or semiring). An infinite ringinal can be thought of as a formalisation of the Leibnizian infinitum terminatum.
AR attempt to argue from “first philosophical principles” that an infinite quantity would necessarily rely on the existence of an infinite whole, thereby contradicting a basic premise in Leibniz. However, the result just mentioned undermines their argument in a mathematically demonstrable way: infinite ringinals are compatible with Euclid’s part-whole axiom, and do not require infinite numbers à la Cantor.
3.4. “Bias against NSA”?
Historians who received their training in mathematics in the traditional Weierstrassian paradigm based on naive Cantorian set theory appear to have difficulty understanding the distinction between unlimited integers, on the one hand, and Cantorian infinities (ordinals and cardinals), on the other. To give an example, AR misunderstood the following passage in Katz et al. [34]:
[One] notes that Leibniz stresses the distinction between infinite cardinality and infinite quantity (reciprocal of infinitesimals) [and] argues that the exchange with Bernoulli precisely refutes [the] attempt to blend infinite cardinality and infinite quantity so as to deduce the inconsistency of infinitesimals. [^37] Katz et al. [34, p. 264].
AR have misread the above passage entirely. They write:
Katz, Kuhlemann, Sherry, Ugaglia and van Atten (2021) maintain, for example, that Leibniz’s argument against Bernoulli is only an argument against infinitesimals interpreted as reciprocals of infinite multiplicities; if there is no number corresponding to an infinite multiplicity, then there can be no reciprocal of such a number either. [^38] Arthur and Rabouin [6, note 27].
So far, AR’s summary is accurate. They continue as follows, still referring to [34]:
This, they claim [i.e., Katz et al. claim], leaves open the question of continuous quantities (magnitudes), and infinitesimals as their reciprocals. Notice that this argument is clearly an invention made in order to save their interpretation from criticism. Until 2016, Katz and co-authors regularly presented infinitesimal magnitudes as a kind of numbers (see the conclusion of Bascelli et al. 2016, 861) and never mentioned the distinction between numbers and magnitudes as a crucial one (see, e.g., Bair et al., 2013, 896). [^39] Ibid.; emphasis on ‘they claim’, ‘continuous’, ‘an invention’, and ‘numbers and magnitudes’ added.
However, the authors of [34] claimed no such thing. The distinction is not between discrete and continuous quantity or magnitude, contrary to what is claimed by AR. The distinction is not between number and magnitude either, contrary to what they claim. Rather, the distinction is the Leibnizian distinction between magnitude and multitude; [^40] For details see e.g., [33, Section 4.1]. in other words, between magnitude or quantity (as formalized by unlimited numbers in modern theories e.g., PA, as discussed in Section 3.3), on the one hand, and infinite wholes formalized as Cantorian infinities (whether ordinal or cardinal), on the other. [^41] Leibniz has a more general notion of an ‘infinite whole’ that includes unbounded lines and unbounded regions. See the beginning of Section 3 and sources given there for a discussion of the paradox of unbounded regions in an exchange of letters with Bernoulli. If there is any invention involved, it is due to AR rather than the authors of [34]. Such a misunderstanding and an attendant conflation are a persistent feature of AR’s writing on Leibniz, as when they write:
[Leibniz] stumbles upon the impossibility of infinite wholes as early as 1672, although he explicitly rejects actual infinitesimals only in 1676.424242Arthur and Rabouin [6, p. 36]; emphasis added.
But infinite wholes are multitudes, whereas infinitesimals are magnitudes. The former are not merely impossible but actually contradictory according to Leibniz; the latter are not contradictory.
In this connection, we would like to comment on AR’s claim that
[T]he matter has often taken a controversial twist after Robinson, any objection to the faithfulness of this reading being seen as relying on an ideological bias against NSA in its various forms. [^43] Op. cit., p. 27; emphasis added.
Recent scholarship (e.g., [8]) has argued that the Leibnizian concept of infinitum terminatum constitutes a non-Archimedean phenomenon, which can be formalized by unlimited numbers in modern NSA or in weaker theories of Peano Arithmetic. The issue is not allegations of “bias against NSA” (as AR claim) but rather the fact that non-Archimedean conceptions are not sufficiently taken into account by AR in their analysis of Leibniz’s mature views. AR announce a follow-up in [7]; it remains to be seen whether it is more coherent than their [6].
4. Homogeneity and comparability
AR present Leibniz’s definition of number from GM VII, 31:
Number is homogeneous with unity, and so it can be compared with unity by adding to it or subtracting from it. And it is either an aggregate of unities which is called an integer, … or an aggregate of aliquot parts of unity, which is called a fraction…” [^44] Leibniz as translated in op. cit., p. 30.
They conclude:
But this definition of number is incompatible with treating an infinitesimal as a number in relation to 1 and “infinitesimals” are never included by Leibniz in his lists of numbers … Infinitesimals cannot be numbers, according to Leibniz, since they are not homogeneous with 1. They cannot be “compared with unity.” [^45] Op. cit., p. 31; emphasis in the original.
Here AR took the Leibnizian remarks out of their context. Leibniz’s definition of number does contain the term ‘compared’, but in the sense of being able add and subtract them (i.e., of being homogeneous), not in the sense of Euclid’s Definition V.4.464646According to Euclid Book V, Definition 4, two homogeneous quantities are comparable if one, multiplied by a finite number, can exceed the other. Ordinary and infinitesimal numbers can indeed be added and subtracted together, e.g., $2+dx$ or $2-dx$.
Note that this Leibnizian definition mentions both number and fraction. By AR’s logic, fractions would similarly be incompatible with infinitesimals. If so, how would AR account for Leibniz’s definition of infinitesimal as a fraction, namely “[an] infinitely small fraction, or one whose denominator is an infinite number”474747“fraction infiniment petite, ou dont le denominateur soit un nombre infini” Leibniz [47, p. 93]. ?
4.1. Leibniz on Euclid V.4 in 1695
Regrettably, AR keep ignoring the fact that in 1695 Leibniz made it clear in two separate texts that infinitesimals violate Euclid’s Definition V.4.484848Leibniz [43, p. 288], [44, p. 322]. For a detailed analysis see [8, Section 1.5, pp. 177–179]. Neither of these Leibnizian texts is even cited in [6], even though AR appear to acknowledge that the presence of inassignables would signal non-Archimedean behavior; see Section 5.4. Instead, they offer non-sequiturs, such as the following:
In the framework of the geometry of the indivisibles, as in the framework of differential calculus, Leibniz’s view is very clear: an infinitesimal $dx$ stands not for a fixed infinitely small quantity but for a finite variable quantity, which one can take as small as one wishes (and the same will be true for infinitely large quantities considered as variable finite quantities taken as great as one wishes). [^49] Arthur and Rabouin [6, p. 35]; emphasis added.
In reality, such an interpretation is far from being ‘very clear’. AR’s claim is merely a restatement of Ishiguro’s position, which remains as unfounded as it was in 1990, as we discuss in Section 4.2. In their Conclusion, AR claim further:
Ever since the publication of the Leibnizian differential algorithm, controversies have occurred amongst mathematicians to determine whether or not it commits its users to accept what are now called non-Archimedean quantities. Leibniz’s own view on that matter, often repeated, was that his algorithm was neutral with regard to these questions. [^50] Op. cit., p. 40; emphasis added.
But Leibniz’s view was not “neutral with regard to these questions” as he made it clear in a pair of 1695 texts that infinitesimals violate Euclid’s Definition V.4.515151See note 48 for the references. Leibniz believed that Archimedean paraphrases are possible, but they amount merely to a reassuring alternative; see further in Section 8.
4.2. Ishiguro on referring and designating
The syncategorematic interpretation of Leibnizian infinitesimals goes back to Ishiguro [28, Chapter 5]. It is the idea that talk about infinitesimals is just stenography for a more laborious argument in a purely Archimedean context, à la exhaustion arguments. The proponents of Ishiguro’s interpretation fiercely reject the idea that mature Leibniz had genuine infinitesimals in mind (i.e., magnitudes in a non-Archimedean context).
Both sides (let’s name them Alice and Bob following [34]) agree that on Leibniz’s mature view, infinitesimals were fictions. However, Alice argues that fiction means stenography as above, whereas Bob argues that fiction means that, unlike his allies l’Hospital and Bernoulli, Leibniz was not convinced that such magnitudes could be found in the physical realm. Being a philosopher, Ishiguro emphasizes the idea that the term infinitesimal does not “refer”:
Fictions … are not entities to which we refer. They are not abstract entities. They are correlates of ways of speaking which can be reduced to talk about more standard kinds of entities. Reference to mathematical fictions can be paraphrased into talk about standard mathematical entities. [^52] Ishiguro [28, p. 100]; emphasis added. Moreover, “The word ‘infinitesimal’ does not designate a special kind of magnitude. It does not designate at all” (Op. cit., p. 83).
Since such a use of the verb refer may not be familiar to our reader, we note that “correlates of ways of speaking” means that talk about infinitesimals is shorthand/stenography for more laborious arguments, etc. A detailed rebuttal of Ishiguro’s reading appeared in [11]. For our purposes, it will suffice to examine Ishiguro’s purported source for the syncategorematic interpretation, mentioned early in her Chapter 5:
Leibniz … maintained that one can have a rigorous language of infinity and infinitesimal while at the same time considering these expressions as being syncategorematic (in the sense of the Scholastics), i.e., regarding the words as not designating entities but as being well defined in the proposition in which they occur. (Letter to Varignon, 2 February 1702, G. Math. 4 93). [^53] Ishiguro [28, p. 82]; emphasis added.
The parenthetical reference is to Leibniz [47, p. 93]. In her Chapter 5, Ishiguro seeks to create the impression that the letter to Varignon provides evidence for her interpretation. Does Leibniz apply the term syncategorematic to infinite magnitudes in this letter? We will examine the issue in Section 4.5.
4.3. Conflation of physical and mathematical realms
Arthur has a history of conflating Leibniz’s physical and mathematical realms, and applying actuality to the mathematical realm. Thus, he claims the following against three Leibniz scholars:
(contra Breger, Bosinelli and Antognazza) [^54] The reference is to Antognazza [1]. there is not one kind of infinity that applies to mathematical entities, and another that applies to actual things. [^55] Arthur [5, p. 158]; emphasis added.
Furthermore, “Leibniz’s actual infinite, understood syncategorematically, [^56] For an analysis of the syncategorematic interpretation see Section 4.2. applies to any entities that are actually infinite in multitude, whether numbers, actual parts of matter, or monads.” [^57] Op. cit., Abstract. There are additional passages of this sort: “Leibniz’s actual infinite, understood syncategorematically, applies to numbers and to constituents of the actual world in exactly the same way” (Op. cit., p. 157); “the actual infinite pertaining to the created world, insofar as it is an infinite in multitude, is syncategorematic in exactly the same sense as the actual infinite that applies to numbers” (Op. cit., p. 158).
After quoting a Leibnizian passage on the actually infinite division of bodies, Arthur comments: “This is exactly the same notion of the actual infinite in multitude as occurs in his mathematics.” [^58] Op. cit., p. 166. But bodies are not the subject matter of mathematics.
On the next page, we are relieved to learn that “Antognazza … insists that the actual infinite cannot apply to ideal entities like numbers” [^59] Op. cit., p. 167, note 16. and we concur with Antognazza.
By page 167 (twelve pages into his article), we locate the source of Arthur’s syncategorematic actual interpretation for numbers, namely Ishiguro:
Thus as Hidé Ishiguro has affirmed, for Leibniz “there are infinitely many substances and infinitely many numbers in his sense of infinitely many, i.e., actually more than any finite number of them, and not merely potentially more” (Ishiguro 1990, 80). [^60] Op. cit., p. 167; emphasis added.
Attached to this sentence on page 80 in Ishiguro’s text is a footnote 2. The footnote provides her two sources. The sources are passages from (a) Leibniz’s letter to Foucher from 1692/93, and (b) Leibniz’s New Essays. The passages deal with division of ‘matter’, ‘creatures’, and ‘things’. They say nothing about ‘numbers’ and therefore don’t support her (and subsequently Arthur’s) application of actuality to numbers. Does Arthur provide any primary sources for his application of ‘syncategorematic actuality’ to numbers? We examine the issue in Section 4.4.
Arthur writes furthermore that “the syncategorematic infinitesimal is not merely an indeterminate finite part: it is a fictional actually infinitely small element, whose use may be justified by exploiting the Archimedean property of the continuum, that is, by substituting for it sufficiently many sufficiently small parts to demonstrate that the error is null.” [^61] Op. cit., p. 170; emphasis added.
Finally, “The actually infinite division of actual matter, syncategorematically understood, issues in infinitely small parts that are not so small that there are not smaller (finite) parts.” [^62] Op. cit., p. 177; emphasis added.
In conclusion, Arthur’s talk about syncategorematic actual infinity for numbers and infinitesimals, following Ishiguro, represents a conflation of the physical and the mathematical realms, is unsupported by evidence, and must be rejected, as already indicated by Antognazza (whose opinion Arthur quotes but fails to follow).
4.4. Arthur’s primary sources
In an attempt to provide a primary source for his application of ‘syncategorematic actuality’ to numbers, Arthur presents Leibniz’s New Essays to the effect that:
“It is perfectly correct to say that there is an infinity of things, i.e. that there are always more of them than can be specified. But it is easy to demonstrate that there is no infinite number, nor any infinite line or other infinite quantity, if these are taken to be genuine wholes. The Scholastics were taking this view, or should have been doing so, when they allowed a syncategorematic infinite, as they called it, and not a categorematic one.” (A VI 6, 157/NE 157). [^63] Leibniz as reported in Arthur [5, p. 161].
Thus far, Leibniz. Arthur then proceeds to claim the following:
This is exactly the same notion of the actual infinite in multitude as occurs in his mathematics: the number of parts (or bodies) is actually infinite in the sense that for any finite number, there are actually (not merely potentially) more bodies than this. It is also precisely the wording Leibniz had used in the New Essays to describe what the Scholastics meant, or ‘should have’ meant ‘when they allowed a syncategorematic infinite’: ‘an infinity of things, i.e. that there are always more of them than can be specified’.  [^64] Arthur [5, p. 166].
Arthur concludes:
I take the implicit criticism in the ‘should have meant’ to be that many Scholastics had followed Aristotle in identifying the syncategorematic infinite with the potential infinite, with the result that they denied the actual infinite. [^65] Ibid.; emphasis added.
What was the ‘implicit criticism’ that Arthur attributes to Leibniz? We gather that it is not as though New Essays ever mentioned ‘syncategorematic actuality’ for numbers; nonetheless, Arthur speculates that this is what Leibniz must have had in mind when he wrote that the Scholastics ‘should have meant’. Such evidence is clearly inconclusive.
There is a brief “supplementary study” [49] where Leibniz seems to illustrate syncategorematic actuality by numbers, but it turns out that the study was “Written on a separate slip of paper and crossed through by Leibniz” [53, p. 409]; furthermore, the supplementary study was never sent to des Bosses. It seems ill-advised to use such a flimsy study as a basis so as to re-evaluate (as Arthur attempts to do in [5, Section 10.5]) Leibniz’s well-established position concerning actuality as applicable only to actual things.
4.5. Analysis of letter to Varignon
As noted in Section 4.2, Ishiguro seeks to present Leibniz’s mention of the syncategorematic infinite as evidence for her interpretation. Let us consider the Leibnizian passage in the letter to Varignon mentioning the syncategorematic infinite. Leibniz wrote:
Yet we must not imagine that this explanation debases the science of the infinite and reduces it to fictions, for there always remains a ‘syncategorematic’ infinite, as the Scholastics say … [^66] Leibniz [47, p. 93] as translated by Loemker (p. 543).
Does Leibniz mention the syncategorematic infinite in reference to an infinite multitude or an infinite magnitude (see Section 3.4)? It suffices to consider the second clause of Leibniz’s sentence:
And it remains true … that $2=\frac{1}{1}+\frac{1}{2}+\frac{1}{4}+\frac{1}{8}+\frac{1}{16}+\frac{1}{32}+\,\cdots$, which is an infinite series containing all the fractions whose numerators are $1$ and whose denominators are a geometric progression of powers of $2$, although only ordinary numbers are used, and no infinitely small fraction, or one whose denominator is an infinite number, ever occurs in it. [^67] Ibid., as translated by Loemker (pp. 543–544).
Leibniz points out that even though the multitude of the terms of the series is infinite, it remains true that no infinite fractions or infinite denominators are involved in summing the series. Thus, the term syncategorematic in this passage does not refer to an infinite magnitude.
In modern mathematics, an infinite multitude is formalized e.g., by a Cantorian cardinal, [^68] Leibniz has a more general notion of an ‘infinite whole’; see note 41. whereas an infinite magnitude is formalized by a nonstandard integer (see Section 3.3). In his letter to Varignon, Leibniz proceeds to discuss the multitude (not magnitude) of terms in a geometric series. Leibniz would of course reject the idea that such a multitude exists as an ‘infinite whole’, and would view it, instead, syncategorematically (one can take as many of the terms as one wishes).
Here Leibniz reiterates the point he made in his correspondence with Bernoulli four years earlier, namely that convergent infinite series cannot be used as evidence for the existence of infinitesimals. The passage does not support the inference sought by Ishiguro. [^69] For further analysis of Ishiguro’s hypothesis, see [11, Section 7.1].
According to Arthur,
Ishiguro claims that Leibniz ‘maintained that one can have a rigorous language of infinity and infinitesimal while at the same time considering these expressions as syncategorematic (in the sense of the Scholastics), i.e., regarding the words as not designating entities but as being well defined in the proposition in which they occur’ (1990, 82), citing Leibniz’s words to this effect in his letter to Varignon of 1702.707070Arthur [5, p. 168, note 19]; emphasis added.
Arthur claims that Leibniz’s letter to Varignon corroborates Ishiguro’s hypothesis. However, as we showed in this section, Leibniz’s words in his letter to Varignon were not at all to Ishiguro’s “effect”.
5. Arthur and Rabouin’s erroneous reading of Robinson
Thus far, we have not mentioned either Abraham Robinson or his theory NSA, except via direct quotations from Arthur and Rabouin. The reason is that AR’s argument is more sweeping than merely rejecting NSA as a historiographic tool; they reject any attribution of genuine infinitesimals to the Leibnizian calculus (and consequently criticize the work of Bos; see further in Section 9), making the question of the applicability of NSA a secondary issue.
However, it would be remiss not to address AR’s misinterpretation of Robinson’s discussion concerning the Leibnizian principle “rules of the finite succeed in the infinite, etc.” as mentioned in their note 4.717171Arthur and Rabouin [6, p. 27, note 4]. We analyze the matter in Section 5.1.
5.1. When do the rules of the finite succeed?
In their comments on Robinson’s historical chapter, AR focus on the issue concerning the precise nature of the “principle” that Leibniz is referring to, and claim that Robinson misidentified the principle. Regardless of the merits of their case, this issue is of secondary importance since the main thrust of Robinson’s comment lies elsewhere: namely, the analogy with the transfer principle. The Leibnizian principle “rules of the finite succeed in the infinite” was not at all obvious to mathematicians at the time, and in fact was the main point of Leibniz’s disagreement with Nieuwentijt. Namely, Nieuwentijt claimed that the square of an infinitesimal is precisely zero; see e.g., Vermij [63]. Nieuwentijt’s stance evidently contradicts Leibniz’s comment concerning the rules of the finite (and similarly, it would contradict the transfer principle).
AR’s footnote 4 starts with a quotation from Robinson:
“The same principle [i.e. Principle of Sufficient Reason] [^72] The bracketed comment was added to Robinson’s passage by Arthur and Rabouin. is supposed to serve as justification for the assumption that the infinitely small and infinitely large numbers obey the same laws as ordinary (real) numbers. Elsewhere, Leibniz appeals to his principle of continuity to justify this assumption.”
This passage appears on page 263 in Robinson [59]. AR go on to comment on Robinson’s passage on page 263 as follows:
[Robinson’s] first sentence rests on a misunderstanding of the letter to Varignon from February 1702 in which Leibniz claims that the “rules of the finite succeeds [sic] in the infinite as if there were atoms” and that, “vice versa, the rules of the infinite succeeds [sic] in the finite as if there were metaphysical infinitely smalls” (although it is not the case that there are such entities in reality). He then concludes: “c’est parce que tout se gouverne par raison”. But here, as in many other similar passages, the recourse to the principle of reason is not a way to justify the transfer between finite and infinite, but to explain why there are no atoms or true infinitesimals in reality. [^73] Arthur and Rabouin [6, note 4].
Here AR miss the point. Robinson argued that his transfer principle can be thought of as a formalisation of heuristic principles found in Leibniz, such as his idea that “the rules of the finite succeed in the infinite … and vice versa, the rules of the infinite succeed in the finite, etc.” as we elaborate in Section 5.2.
5.2. Transfer principle
To give an elementary illustration, the transfer principle entails that a formula true for all real inputs would necessarily remain true also for all hyperreal inputs. For instance, consider the formula expressing the idea that if a real number is nonzero, then its square will also be nonzero, or in symbols
$\forall x\in{\mathbb{R}}\,(x\not=0\to x^{2}\not=0).$
By transfer, the square of any nonzero infinitesimal must similarly be nonzero. To give another example, consider the trigonometric identity $\sin^{2}x+\cos^{2}x=1$ for all $x\in{\mathbb{R}}$. The trigonometric functions possessing natural extensions to the hyperreal domain, the transfer principle would therefore assert that the formula remains true for all infinitesimal or infinite $x$ (as well as all other hyperreal $x$).
In their objection, AR miss the point of Robinson’s insight. Which philosophical principle precisely Leibniz referred to in his letter to Varignon is a secondary issue, the main issue being the parallelism between Leibniz’s idea of finite rules succeeding in the infinite, on the one hand, and Robinson’s transfer principle, on the other. AR’s claim that Robinson misunderstood the letter to Varignon is based on their own misunderstanding of the main thrust Robinson’s comment on page 263. Robinson is even clearer on page 266:
Leibniz did say, in one of the passages quoted above, that what succeeds for the finite numbers succeeds also for the infinite numbers and vice versa, and this is remarkably close to our transfer of statements from ${\mathbb{R}}$ to ${{}^{\ast}{\mathbb{R}}}$ and in the opposite direction. [^74] Robinson [59, p. 266].
Leibniz’s heuristic principle (the Law of Continuity) was not a self-evidence at the time, and in fact lay at the heart of his argument against Nieuwentijt’s approach (see [44]). Nieuwentijt postulated a single infinite number, formed an infinitesimal by passing to its reciprocal, and postulated that the square of the infinitesimal is zero. Nieuwentijt’s infinitesimal obviously does not adhere to the rule “whatever succeeds for the finite, succeeds also for the infinite” (namely, it contradicts the rule that a nonzero number must have a nonzero square). Parallels between the procedures of the Leibnizian calculus and those of modern infinitesimal analysis are explored in [32].
5.3. Ehrlich on non-Archimedean theories
In their discussion of “non-Archimedean theories of the continuum in general,” AR claim that “there were already good exemplars before the intervention of Robinson, as Philip Ehrlich has argued in detail.” [^75] Arthur and Rabouin [6, p. 28]; emphasis added.
AR’s claim amounts to both a slight toward Robinson’s theory of infinitesimals which was the first theory truly applicable in calculus and analysis broadly understood (unlike the earlier theories), and a misrepresentation of the position of Ehrlich, who wrote clearly:
After mathematicians had been taught for decades that a consistent theory of the calculus based on infinitesimals was impossible, Abraham Robinson was certainly swimming against the tide when he proved otherwise. [17]
Ehrlich did present a detailed analysis of non-Archimedean theories preceding Robinson’s in [16]. However, such theories failed the Klein–Fraenkel criterion of applicability in broader mathematics, as argued in [31]. More specifically, such theories did not possess a sufficiently strong version of the transfer principle discussed in Section 5.2.
5.4. Leibniz, Nelson, and inassignables
Building upon Robinson’s work in NSA in the 1960s [59], Karel Hrbacek [27] and Edward Nelson [55] developed axiomatic approaches to NSA in the 1970s. AR claim that Nelson’s approach bears resemblance to an early period (1669) when Leibniz endorsed actual infinitesimals:
“[O]ne of us [namely, Arthur] has argued that prior to 1676 Leibniz did treat indivisibles as actual but unassignable elements of the continuum in three more or less distinguishable phases of his thought, in one of which they bear some resemblance to the non-Archimedean infinitesimals of Edward Nelson’s Internal Set Theory; see Arthur (2009).” [^76] Arthur and Rabouin [6, p. 28, note 13].
The reference is to Arthur’s 2009 article, [^77] Arthur [3, pp. 15–16]. where an analogy with Nelson’s work is applied to
“Phase 1 (1669): Unassignable gaps in the continuum.” [^78] Op. cit., p. 39.
This phase in Leibniz’s development was analyzed in detail by Ugaglia and Katz. [^79] Ugaglia and Katz [62, Section 3.2.1]. During this period,
[E]xtension is continuous, and continuity means for Leibniz potentially infinite divisibility, that is syncategorematicity, as it was for Aristotle. The terminology itself of unassignability and infinitesimals is absent in this phase: [^80] Two years later, Leibniz will use inassignabilis in his Theoria motus abstracti [39, p. 265]. This text belongs to Phase II of Leibniz’s development. ‘infinitely small’ just means ‘potentially infinitely divisible.’ [^81] Ugaglia and Katz [62, p. 13].
We will analyze this period in more detail in Section 5.5.
5.5. Drift away from syncategoremata
In this early period, Leibniz endorsed the traditional syncategoremata that he learned from the scholastics, which resulted in a mathematically inconsistent approach. There are no unassignable (or inassignable) infinitesimals during this period. Thus Arthur’s comparison of Leibniz’s writings from 1669 and Nelson’s theory is unfounded. [^82] Arthur seems to be aware of this difficulty when he writes “The textual basis for Leibniz’s first theory of the continuum is slender, and has to be pieced together from unpublished fragments and his own later testimony” [3, p. 14]. We argue that Arthur pieced it together incorrectly. Beeley writes:
Having little background in mathematics at that time [i.e., in 1669–71], Leibniz had evidently gathered all he knew about Cavalieri and Wallis and the criticisms which had been directed against their respective methods from the polemical writings of Thomas Hobbes. [^83] Beeley [12, pp. 44–45].
As Leibniz’s knowledge of mathematics grew, Leibniz abandoned the earlier approaches (including syncategoremata) in favor of his fictionalist approach to genuine infinitesimals.
It is instructive to ponder the implications of AR’s discussion of Nelson in connection with Leibniz’s alleged period of “Unassignable gaps in the continuum.” Such discussions, found in both Arthur (2009) and Arthur and Rabouin (2024), appear to acknowledge that the presence of unassignables (or inassignables) would signal non-Archimedean behavior that “bears some resemblance” to Nelson’s theory, as already mentioned in Section 4. AR’s insights here could be usefully extended to understanding Leibniz’s mature views – a route regrettably not taken by AR.  Curiously, Arthur does envision comparing Leibniz’s mature views with a modern theory called Smooth Infinitesimal Analysis (SIA) in [4] (with the unsurprising conclusion that they are not similar because SIA has nilpotent infinitesimals); see [8] for an analysis.
6. Hornangles vs inassignables
Hornangles (also known as angles of contingence, or angles of contact) were discussed in Euclid and debated in the 16–17th centuries; see Maierù [54]. AR’s discussion of hornangles overlooks the fact that they are a manifestation of non-Archimedean behavior, as acknowledged e.g., by Sonar, who writes:
[A]lready Eudoxus knew that also other number systems – so-called non-Archimedean number systems – were conceivable … Such a system of quantities which was already known to the Greeks were cornicular angles or horn angles. [^84] Sonar [60, p. 40].
Surprisingly, AR open their section 7 on hornangles with yet another conflation of non-Archimedean phenomena and ‘nonstandard quantities’, as follows: “it might be useful to analyse a last example regularly mentioned in support of the existence of nonstandard quantities in Leibniz: horn angles.” [^85] Arthur and Rabouin [6, p. 39]; emphasis added. AR fail to provide any quotation of an alleged attribution of a “nonstandard” attribute to hornangles, and with good reason: we are not aware of any such attribution. [^86] The reference [33, Section 4.5] cited by AR in footnote 99 on page 39 of [6] does contain a discussion of hornangles, but certainly makes no such attribution, contrary to AR’s claim.
6.1. Euclidis Prôta: page 191
AR’s main source on hornangles is Leibniz’s 1712 work Euclidis Prôta [50]. They present the following passage on page 191 in GM:
“Indeed, it is not possible to assign a line which would be to another line as is the angle of contingency $DCB$ at the angle of contingency $DCF$, which we already showed to be possible for rectilinear angles.” [^87] Leibniz as translated by Arthur and Rabouin [6, p. 39, note 101].
AR go on to translate an additional passage from page 191:
“And therefore, they are from a completely different genre, and in relation to rectilinear angles, it [scil. the angle of contact] [^88] The bracketed comment is in AR’s text. cannot even be considered as an infinitely small, which is always located between zero and an assignable.” [^89] Leibniz as translated by Arthur and Rabouin [6, p. 40]. The original reads as follows: “Ac proinde plane est diversi generis, et respectu anguli rectilinei ne quidem ut infinite parvus considerari potest, qui utique inter nullum et assignabilem collocatur” [23], vol. V, p. 191.
AR proceed to contest the conclusions of the analysis of Leibniz’s use of hornangles as developed by Katz et al. [33, Section 4.5]. AR rely mainly on these two quotations from page 191 in GM to develop their rebuttal. While they mention the year, 1712, of the Leibnizian text Euclidis Prôta, they omit to mention the year of the Leibnizian text that provided the basis for the analysis in [33]. That year is 1686 (Leibniz [42]), namely no fewer than 26 years prior to Euclidis Prôta.
6.2. Leibniz on Peletier and Clavius
AR similarly fail to quote a remarkable passage from page 191 in Euclidis Prôta that clarifies Leibniz’s position on both hornangles and inassignables. It turns out that Leibniz sides with Peletier in his debate with Clavius (who endorsed hornangles), declares that Euclid didn’t really mean it when he described hornangles as quantities, and says that Clavius’ endorsement of hornangles was a great embarrassment for geometers, and provided fodder for Hobbes’s ridicule of geometers:
“And so in this matter I agree with Peletarius against Clavius; and Euclid, when he said that the angle of contact is smaller than any straight line, spoke a little too loosely, understanding “smaller” to mean something whose beginnings fall within the space of the former. But it should not therefore be considered that he has attributed a perfect quantity to the angle of contact with respect to the straight line. … and Clavius erred greatly, when he denied this axiom, by which it is asserted that what passes from one extreme to another, must also equally pass through everything in between, and on this account gave Thomas Hobbes an opportunity to insult geometers.” [^90] Leibniz [50, p. 191]; translation ours.
One is struck by the contrast between enthusiastic bashing of hornangles in 1712, on the one hand, and their enthusiastic endorsement 26 years earlier in his study of the degrees of contact among curves in [42], on the other. The contrast indicates that in the meantime Leibniz had had a change of heart concerning hornangles.
6.3. An infinitely small between zero and an assignable
But apart from the issue of hornangles, the most pertinent fact about page 191 of Euclidis Prôta is that it contains a sentence that directly refutes AR’s insistence that Leibnizian infinitesimals do not testify to a non-Archimedean phenomenon. In fact, it is a passage that was quoted by AR, who appear to have overlooked its implications. Indeed, when Leibniz writes that
“an infinitely small is always located between zero and an assignable,” [^91] See main text at note 89.
he reveals that an infinitely small is inassignable and therefore violates Euclid’s Definition V.4, as Leibniz made clear elsewhere. [^92] See note 48 for the references.
A hornangle, while being smaller than every assignable rectilinear angle, is not accorded the status of a genuine quantity by Leibniz in 1712. Unlike a hornangle, an infinitesimal, similarly smaller than every assignable quantity, is indeed accorded the status of quantity as per page 191 of Euclidis Prôta.
There is a fundamental incoherence in AR’s position with regard to inassignables (or unassignables). If inassignables are contradictory and/or are to be understood syncategorematically (see Section 4.2), why does Leibniz keep mentioning inassignables, as for example when he describes the characteristic triangle as being of inassignable size in his letter to Wallis, as analyzed by Beeley? [^93] Leibniz [46], quoted in Beeley [12, p. 48, note 36]. Beeley further mentions that “[Leibniz], in a letter to Wallis, … sets out reasons why the geometry of indivisibles cannot strictly be reduced to the ancient method of exhaustions, noting that the one operates with finite quantities, the other with quantities incomparably smaller than the whole” [12, p. 51]. Here Leibniz states clearly that his method does not operate merely with finite quantities, contrary to Ishiguro’s interpretation erdorsed by AR. Leibniz specifically objects to Wallis’s habit of describing infinitesimals as being exactly zero. Leibniz points out that such an approach introduces unnecessary obscurity into analysis. Arguably, dealing in contradictory entities would similarly introduce unnecessary obscurity, as would be describing a syncategorematically-inspired sequence of ordinary, assignable values getting smaller and smaller as being inassignable. [^94] See main text in Section 7.2 at note 97 for a related point concerning bounded infinities.
7. Proposition 11 and infinita terminata
In his De Quadratura Arithmetica, Leibniz introduces the notion of infinita terminata (literally “bounded infinites”) and contrasts it with unbounded infinites, or infinite wholes; see Section 3.
7.1. Should Leibniz be taken literally?
In this connection, AR write:
Leibniz’s strategy in prop. XI is to shortcut any reference to this ‘unbounded’ space (for which the notion of area would be problematic) by introducing a parallel line to the asymptote which can be taken as close as one wishes but will still intersect the given curve. … But that means that the line, which was introduced to perform the quadrature, can be made as long as one wants (greater that any given magnitude). [^95] Arthur and Rabouin [6, p. 34].
They go on to the literal meaning:
Taken literally (as a fixed and not as a variable object), it hence describes a line which is at the same time bounded (it intersects the curve at a point and we can make it begin at the origin of our frame of coordinates) and infinite (it can be made larger than any given quantity). [^96] Op. cit., pp. 34–35; emphasis added.
Note that AR are not quoting Leibniz here but only paraphrasing him. How do they mean exactly to say that “Taken literally (as a fixed and not as a variable object), it hence describes a line which is at the same time bounded … and infinite”? Did Leibniz mean it literally, or did he mean to say that “it can be made larger than any given quantity” as AR claim? These are apparently two different things, in spite of AR’s effort to blur the distinction. Either Leibniz meant it literally, or else what he had in mind was a paraphrase along the lines “can be made as long as one wants”.
It is significant that AR don’t quote Leibniz directly, because if one examines what Leibniz actually wrote (see for instance [35]), there is little doubt as to what he meant, as we discuss in Section 7.2.
7.2. Leibniz’s infinitesimal $(\mu)\mu$
In Proposition 11, Leibniz proceeds as follows.
  1. (1)
    He denotes a point at infinitesimal distance from the origin by $(\mu)$, where the parentheses are part of the notation (whereas $\mu$ itself denotes the origin).
  2. (2)
    He considers a line parallel to the asymptote and at infinitesimal distance from it (i.e., passing through the point $(\mu)$ as above).
  3. (3)
    He takes the corresponding point on the curve, whose ordinate is then a specific fixed (and not a variable) infinite entity, namely an infinitum terminatum.
AR’s paraphrase of Leibniz rests upon the assumption that he must have meant it syncategorematically (see Section 4.2). In doing so, they create an impression that a genuine bounded infinity would have to be as contradictory as a quantity that is both fixed and variable.
AR’s conclusion with regard to the Leibnizian bounded infinity is that “the line, which was introduced to perform the quadrature, can be made as long as one wants (greater that any given magnitude)” as quoted above. However, this begs the question concerning their Ishiguro/exhaustion reading. According to their explanation, there is no difference between bounded infinity and an (arbitrarily large) finite line segment. If so, what would be the point for Leibniz to introduce a paradoxical-sounding term (“bounded infinity”!) if he simply meant ‘finite line segment’? This amounts to a basic incoherence of their intepretation. [^97] See main text in Section 6.3 at note 94 for a related point concerning infinitesimals.
7.3. Three approaches to bounded infinity
How does Leibniz deal with the contradictory nature of an infinite whole such as an unbounded line? He dissolves the contradiction by saying that one can take a finite line bigger than any given, i.e., by viewing infinity syncategorematically (not as a whole). How do AR interpret bounded infinity? Also by saying that one can take a finite line bigger than any given.
It emerges that, on AR’s reading, infinitum terminatum is indistinguishable from unbounded infinity. Such an interpretation however is incoherent because Leibniz explicitly contrasts them in the Scholium to Proposition 11.989898See note 25.
There are three interpretations of infinitum terminatum in the current literature:
  1. (1)
    As argued in [35], on Knobloch’s reading it amounts to an ideal perspective point at infinity. [^99] Briefly, Knobloch analyzes theorem 45 in Leibniz’s De Quadratura Arithmetica, and claims that the point at infinity introduced there is an instance of the infinitum terminatum. However, that point at infinity is recognizably the ideal perspective point at infinity known already to Kepler and Desargues, as argued in [35].
  2. (2)
    On AR’s Ishiguro-inspired reading, it is indistinguishable from unbounded infinity.
  3. (3)
    We argue that it is an inverted infinitesimal and thus a genuine mathematical entity not found in the physical realm.
The weaknesses of interpretations (1) and (2) are analyzed respectively in [35] and the present text.
8. Question of use and question of existence
In interpreting historical infinitesimalists, it is important to distinguish between mathematical practice, on the one hand, and foundational ontology, on the other. Modern analysis, whether of Weierstrassian or Robinsonian variety, is typically built upon set-theoretic foundations involving infinite sets as a sine qua non (but see Section 3.3). Such foundations are far removed from the world of the historical infinitesimalists such as Leibniz, Euler, and Cauchy. On the other hand, the procedures of modern analysis, whether of Weierstrassian or Robinsonian kind, often find close analogs in the work of historical infinitesimalists.
It is obvious that a study of Leibnizian calculus must focus on the procedures if it is to avoid the trap of presentism. We can only study the procedures Leibniz used in his calculus and geometry, and certainly cannot attribute modern set-theoretic investigations to him. Thus the issue of use should be clearly separated from the issue of existence as a mathematical entity in a set-theoretic framework.
AR appear to set this distinction on its head when they claim that
[A]s long as one conflates questions of use and questions of existence, it will always be possible to counterbalance any explicit denial of the existence of infinitesimals by Leibniz by an appeal to his practice as testifying to the contrary. At a minimum, one could argue that there is tension between what he says and what he does.100100100Arthur and Rabouin [6, p. 27]; emphasis added.
However, this is not at all what has been argued in recent Leibniz scholarship on his genuine infinitesimals. Leibniz clearly denied that infinitesimals exist in the phenomenal realm, and they are therefore fictional entities like negatives and imaginary roots. Contrary to AR’s claim, such a reading does not constitute a “conflation of the question of use with that of existence of infinitely small quantities.” AR go on to claim that
Leibniz himself emphasized on many occasions that his practice should remain neutral as regards the question of existence, and henceforth that the use of infinitesimals should be interpretable in both ways (i.e., either as existing entities or not).101101101Op. cit., p. 28.
Their parenthetical comment reveals a basic misunderstanding of the Leibnizian position. On many occasions, Leibniz emphasized that the use of infinitesimals can be replaced by equivalent (but more laborious) procedures à la Archimedean exhaustion. It is misleading to characterize the existence of two distinct methods as treating infinitesimals “either as existing entities or not” since by all accounts, they do not exist in the phenomenal realm, according to Leibniz. The issue is whether or not to use them directly. AR conclude that
[One] needs to explain how it is that one can employ fictions of this type (contradictory notions) in mathematical discourse without falling into inconsistency. Moreover, if one conflates questions of use and questions of existence, this explanation will not alter the impression that there is a tension between Leibniz’s justification and his practice.102102102Op. cit., p. 41; emphasis added.
If there is any ‘tension’, it is created by AR’s misconceived idea that infinitesimals are ‘contradictory notions’. The alleged conflation between use and existence is a made-up problem, as explained above.
9. Two syncategorematists against Bos
The influential study of Leibnizian methodology by Bos [14] was published in 1974. Both Arthur and Rabouin are on record as criticizing the study by Bos whenever his analysis creates tensions with Ishiguro’s 1990 reading (see Section 4.2).
In 2020, Rabouin and Arthur quote an example of Leibniz’s application of the Law of Continuity, and then proceed to criticize Bos in the following terms:
This first basic example was not mentioned in Henk Bos’s seminal study (Bos 1974–5), and this had the unfortunate consequence of hiding the global structure of the reasoning.103103103Rabouin and Arthur [58, p. 439]; emphasis added.
The alleged ‘global structure’ is Ishiguro’s syncategorematic reading. In reference to an additional example, they lodge a further criticism:
Here again, the presentation by Bos, as thorough and interesting as it is, had the unfortunate consequence of hiding some particularities of the reasoning by neglecting some developments that were apparently of no interest.104104104Ibid.
Related criticisms by Arthur appear in an earlier text dating from 2013.
9.1. Arthur vs Bos, 2013
In 2013, Arthur claimed the following:
[I]t is not the case that Leibniz has two methods, one committed to the existence of infinitesimals as ‘genuine mathematical entities’, ‘fixed, but infinitely small’, and the other Archimedean, treating the infinitely small as arbitrarily small finite lines, ones that could be made as small as desired.105105105Arthur [4, p. 561]; emphasis added.
Bos’s position was precisely that Leibniz did have two methods; we concur. Arthur continues criticizing Bos in footnote 5 on the same page:
Bos, in his classic article on Leibniz’s differentials (1974–1975, 55), sees Leibniz as pursuing “two different approaches to the foundations of the calculus; one connected with the classical methods of proof by ‘exhaustion’, the other in connection with a law of continuity.” The quotations given in my text are from Bos (1974–1975), 12, 13. See also Herbert Breger’s similar criticisms of Bos on these points in his (2008, 195–197).106106106Op. cit., note 5; emphasis added.
Arthur goes on to contrast his work with that of Bos:
My interpretation here contrasts with that of Bos, who reads Leibniz as having two different interpretations of differentials, one in which they are infinitely small elements of lines, and the other where they stand for finite lines.107107107Op. cit., p. 566.
Here Bos summarized Leibniz’s approach of using $dx,dy$ for infinitesimal differentials, and a separate notation $(d)x,(d)y$ when “ordinary” values are substituted for $dx,dy$108108108In his 1974 article, Bos used $\underline{d}x$ in place of Leibniz’s $(d)x$. After outlining the reading by Bos, Arthur goes on to lodge an even stronger claim of “misinterpretation”:
But this, it seems to me, is to misinterpret Leibniz’s justification, …109109109Arthur [4, p. 567]; emphasis added.
9.2. Who is Arthur and Rabouin’s quarrel with?
It emerges from the analysis above that AR’s quarrel is neither with Robinson, nor Nelson, nor yet “the mathematician Mikhail G. Katz with various co-authors”110110110Arthur and Rabouin [6, note 2]. but rather squarely with the historian Henk Bos, whose work on Leibniz is described by Guicciardini as “a classic in our field” [15, pp. 40–41]. Naturally, other historians have also disagreed with AR’s interpretation. Thus, the 2021 article by Esquisabel and Raffo Quintana explicitly rejects it in the following terms:
  1. (1)
    “[U]nlike the infinite number or the number of all numbers, for Leibniz infinitary concepts do not imply any contradiction, although they may imply paradoxical consequences.” [20, p. 641]
  2. (2)
    “[W]e disagree with the reasons [Rabouin and Arthur] gave for the Leibnizian rejection of the existence of infinitesimals, and in our opinion the texts they refer to in order to support their interpretation are not convincing. Since we argue that Leibniz did not consider the concept of infinitesimal as self-contradictory, we try to provide an alternative conception of impossibility.” [20, p. 620]
Antognazza’s objections to Arthur’s interpretation were analyzed in Section 4.3. Further objections were detailed in Eklund [18].
9.3. Jesseph vs the syncategorematic interpretation
Jesseph sees Leibniz as having two ways of interpreting the calculus:
Rather than postulate an ambiguity in Leibniz’ writings, we can take him at his word: he is convinced that infinitesimal magnitudes are eliminable, and the reason for this conviction is the thoroughly commonsensical belief that the truth of his mathematical results should not depend too heavily upon the resolution of metaphysical problems. In the Leibnizian scheme, true mathematical principles will be found acceptable on any resolution of the metaphysical problems of the infinite. Thus, Leibniz’ concern with matters of rigor leads him to propound a very strong thesis indeed, namely no matter how the symbols “dx” and “dy” are interpreted, the basic procedures of the calculus can be vindicated.111111111Jesseph [29, p. 243].
Jesseph concludes:
Such vindication could take the form of a new science of infinity, or it could be carried out along classical lines, but in either case the new methods will be found completely secure.112112112Ibid.
Accordingly, infinitesimals are not merely an abbreviated way of speaking (as per to Ishiguro) and certainly not contradictory (as per AR), but rather well-founded fictions. In his 1998 text on page 35, Jesseph explains the difference between well-founded and ill-founded fictions, and concludes:
Infinitesimal magnitudes … are nevertheless well-founded fictions precisely because the realm of mathematical objects is structured just as it would be if the infinitesimal existed.113113113Jesseph [30, pp. 35–36].
In these passages, Jesseph clearly distances himself from the syncategorematic interpretation of Leibniz’s infinitesimals. It is therefore surprising that he should have consented to have his name added to the author list of the broadside [2].
To summarize, we list some of the Leibniz scholars who disagree with the reading by Ishiguro, Arthur, and Rabouin: Antognazza, Bos, Eklund, Esquisabel and Quintana, Jesseph.
10. Arthur vs Rabouin: Limit processes in 2013
In 2013, a pair of texts appeared that offer an interesting comparison. One of them, “Leibniz’s syncategorematic infinitesimals” by Arthur, claims that (syncategorematic) infinitesimals involve a ‘limiting process’. Another, by Rabouin, claims that, unlike the Newtonian tradition, there were no ‘limit processes’ in the Leibnizian calculus. Thus, Arthur wrote:
The role of $(d)y$$(d)v$ and $(d)x$ is to be “finite surrogates”: they remain assignable quantities throughout the limiting process as $dx$ becomes arbitrarily small, and since $(d)y$ and $(d)x$ have an assignable ratio equal to $dy:dx$ even in that limit (likewise for $(d)v$ and $(d)x$), differential equations involving terms in $dy$$dv$ and $dx$ to the same order are interpretable under the fiction where these quantities stand for infinitely small quantities.114114114Arthur [4, p. 567]; emphasis added.
Arthur’s is a clear statement to the effect that the syncategorematic interpretation involves a ‘limiting process’. Meanwhile, Rabouin wrote in the same year:
A profound evolution of the concept involved in differential calculus, already emphasized by D’Alembert, was the progressive introduction of a new interpretation of its basic constituents in terms of “limit processes” – an interpretation that is absent from the practice of Leibniz and his first followers, but present in the Newtonian tradition.115115115Rabouin [56, p. 24]; emphasis added. By 2017, the passage was modified by deleting the phrase “absent from the practice of Leibniz” as follows: “A profound evolution of the concept involved in differential calculus, already emphasized by Jean le Rond d’Alembert, was the progressive introduction of a new interpretation of its basic constituents in terms of limit processes–an interpretation that is put forward in the Newtonian tradition. The Leibnizian calculus developed quite rapidly as a mixture of features derived from Leibniz’s and Newton’s techniques, etc.” [57, p. 213]. The 2013 version was the more accurate one.
We concur with Rabouin’s view as expressed in 2013 that in Leibniz’s mature view of the calculus, there were no limit processes – and therefore no syncategorematics as per Arthur 2013.
If the interpretation in terms of “limit processes” is absent from the Leibnizian calculus as per Rabouin 2013, then infinitesimals are not to be interpreted syncategorematically (as per Arthur 2013), and therefore they are entities exhibiting non-Archimedean behavior as per Rabouin 2013. Since there is no trace of any endorsement of an NSA interpretation of Leibniz in Rabouin 2013, it emerges that Rabouin was capable of envisioning non-Archimedean phenomena in Leibniz without endorsing any concomitant Leibniz–Robinson linkage. Somehow the distinction between genuine infinitesimals and those of Robinsonian analysis has been erased by the time of writing Rabouin’s 2024 text with Arthur.
11. Conclusion
We have analyzed the recent publication by Arthur and Rabouin in Historia Mathematica on Leibniz’s infinitesimals, and identified many errors stemming in a number of cases from flawed analyses in their earlier publications. Their thesis is in fact far more radical than what their title and abstract may suggest. Arthur and Rabouin are not merely rejecting interpretations of the Leibnizian calculus based on Abraham Robinson’s theory. Their claim goes beyond the issue of the historiographic applicability of nonstandard analysis, and amounts to a sweeping claim concerning the Leibnizian calculus itself. Their claim consists in rejecting any attribution of genuine infinitesimals to the Leibnizian calculus.
It is no accident that their article fails to cite the seminal study of Leibnizian methodology by Bos [14], described as “a classic in our field” by Guicciardini [15]. Bos did acknowledge that Leibniz used genuine infinitesimals (and there is no trace of syncategoremata in his study). In sum, Arthur and Rabouin’s claim is unviable.
12. Acknowledgments
We are grateful to Ali Enayat, Karel Hrbacek, David Sherry, Monica Ugaglia, and the anonymous referee for helpful suggestions. The influence of Hilton Kramer (1928–2012) is obvious.